Two angles with measures 2x2 + 3x - 5 and x2 + 11x - 7 are supplementary. Find the value of x.

Answer by stanbon(75887)

Two angles with measures 2x2 + 3x - 5 and x2 + 11x - 7 are supplementary. Find the value of x.
-----
Equation:
2x^2+3x-5 + x^2+11x-7 = 180
----
3x^2 + 14x -12 = 180
----
3x^2 + 14x - 192 = 0
Factor:
(x-6)(3x+34) = 0
---
If x = 6
2x^2 + 3x-5 = 2*36+3*6-5 = 85 degrees
x^2 + 11x - 7 = 6^2+11*6-7 = 95 degrees
=============================================
cheers,
Stan H.
========================